A proton moves through a magnetic field at 28.7% of the speed of light. At a location where the field has a magnitude of 0.00679 T and the proton\'s velocity makes an angle of 115° with the field, what is the magnitude of the magnetic force acting on the proton?

___ N
Answer #1

Here F=q*v x B

where q=1.6 x 10^-19 C
v=0.287*(3 x 10^8 m/s)= 8.61 x 10^7 m/s
B= 0.00653T*Sin115°

Solving for F

F= (1.6 x 10^-19 C)*(8.61 x 10^7 m/s)*(0.00679T*Sin115°)

F= 8.48 x 10^-14 N Answer